OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y + 9)2 = 4. What is the center and radius of circle S? Center: (16, -9); Radius: 4 Center: (-16, 9); Radius: 2 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 4

OpenStudy (anonymous):

\[\small\text{Form of a Circle: } (x - h)^{2} + (y - k)^{2} = r^{2} \text{where the center is (-h, -k) and the radius is r.}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is the answer cause i don't understand

OpenStudy (anonymous):

Alright. h = -16 and k = 9. WHat is (-h, -k) given the values?

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

Alright. If you have h = -16,, then -h = -(-16) and -h = 16. Do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ALright. Given that, if k = 9, then -k = ?

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

Yup. Therefore, the center is (16, -9). Now, for the radius, r = radius. \[r^{2} = 4\]\[\sqrt{r^{2}} = \sqrt{4}\]THerefore, r = ?

OpenStudy (anonymous):

16,-9?

OpenStudy (anonymous):

(16, -9) is the CENTER. What is the value of √4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok. THat means that the radius is 2. THerefore, the center is (16, -9) and the radius is 2.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

np :)